# Preconditioning without a preconditioner: faster ridge-regression and Gaussian sampling with randomized block Krylov subspace methods

### v1

@@ 2.2.1 Exact deflation
Traditionally, it has been suggested to take $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ as $[\![\mathbf{A}]\!]_{r}$, the best rank-$r$ approximation to $\mathbf{A}$; i.e. to set $\mathbf{D}$ as the top $r$ eigenvalues of $\mathbf{A}$ and $\mathbf{U}$ the corresponding eigenvectors. In this case Equation 8 is sometimes called a spectral deflation preconditioner. The name arises because the eigenvalues of $\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}$ are
$\underbrace{\theta+\mu,\ldots,\theta+\mu}_{\text{$r$-copies}},\lambda_{r+1}+\mu,\ldots,\lambda_{d}+\mu,$ (10)
and hence, if $\theta\in[\lambda_{d},\lambda_{r+1}]$, then the condition number of $\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}$ is $(\lambda_{r+1}+\mu)/(\lambda_{d}+\mu)$. In other words, the top $r$ eigenvalues of $\mathbf{A}_{\mu}$ are deflated. In combination with Corollary 2.3, this yields the following convergence guarantee.
@@
Let $\mathbf{P}_{\mu}$ be the preconditioner Equation 8 corresponding to $[\![\mathbf{A}]\!]_{r}$, the rank-$r$ truncated SVD of $\mathbf{A}$, and $\theta\in[\lambda_{d},\lambda_{r+1}]$ and let
$\varepsilon_{t}(\mu):=2\exp\left(-\frac{2t}{\sqrt{(\lambda_{r+1}+\mu)/(\lambda_{d}+\mu)}}\right).$
Then the $t$-th PCG iterate (Definition 2.1) corresponding to the preconditioner $\mathbf{P}_{\mu}$ satisfies
$\frac{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{pcg}_{t}(\mu)\|_{\mathbf{A}_{\mu}}}{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}\|_{\mathbf{A}_{\mu}}}\leq\varepsilon_{t}(\mu).$
When $\lambda_{r+1}\ll\lambda_{1}$ and $\mu$ is not too large, then the rate of convergence guaranteed by the bound can be much faster than without preconditioning ($r=0$). [^3] Standard CG also satisfies bounds in terms of $\lambda_{r+1}/\lambda_{d}$, at least in exact arithmetic; see Theorem D.1. Of course, any potential benefits to convergence must be weighted against the cost of constructing the spectral deflation preconditioner, and exact deflation, which requires computing the top eigenvectors of $\mathbf{A}$, can be costly.
@@ 2.2.2 Nyström preconditioning
Techniques from randomized numerical linear algebra allow near-optimal low-rank approximations to $\mathbf{A}$ to be computed very efficiently (Halko et al., 2011; Tropp & Webber, 2023). Intuitively, we might hope that the the corresponding preconditioner works nearly as well as spectral deflation.
In particular, it’s reasonable to take $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ as the eigendecomposition of the Nyström approximation
$\mathbf{A}\langle\mathbf{K}_{s}\rangle:=(\mathbf{A}\mathbf{K}_{s})(\mathbf{K}_{s}^{\mathsf{T}}\mathbf{A}\mathbf{K}_{s})^{\dagger}(\mathbf{K}_{s}^{\mathsf{T}}\mathbf{A}),$ (11)
where $\mathbf{\Omega}\in\mathbb{R}^{d\times\ell}$ is a matrix of independent standard normal random variables and
$\mathbf{K}_{s}:=[\mathbf{\Omega}~{}\mathbf{A}\mathbf{\Omega}~{}\cdots~{}\mathbf{A}^{s-1}\mathbf{\Omega}]\in\mathbb{R}^{d\times(s\ell)}.$ (12)
This variant of the Nyström approximation is among the most powerful randomized low-rank approximation algorithms, and can be implemented using $s$ matrix-loads (Tropp & Webber, 2023).
Nyström-based preconditioning has proven effective in theory and practice, and has been an active area of research in recent years (Martinsson & Tropp, 2020; Frangella et al., 2023; Carson & Daužickaitė, 2024; Díaz et al., 2023; Zhao et al., 2024; Dereziński et al., 2025).
Most related to the content of this paper is the theoretical analysis of (Frangella et al., 2023), which proves that if $s=1$, $\theta=\lambda_{\ell}(\mathbf{A}\langle\mathbf{\Omega}\rangle)$, and the sketching dimension $\ell$ is on the order of the effective dimension
$d_{\textup{eff}}(\mu):=\operatorname{tr}\bigl{(}\mathbf{A}\mathbf{A}_{\mu}^{-1}\bigr{)}=\sum_{i=1}^{d}\frac{\lambda_{i}}{\lambda_{i}+\mu},$ (13)
then Nyström PCG converges in at most $t=O(\log(1/\varepsilon))$ iterations; i.e. independent of any spectral properties of $\mathbf{A}$. Our analysis makes use of the same general techniques as (Frangella et al., 2023), but is applicable when $s>1$ as well as if $\mu=0$. We discuss the theoretical bounds of (Frangella et al., 2023) as well as some possible shortcomings of using bounds based on the effective dimension in Section D.2.
@@ 3 Our approach: augmented block-CG
In this paper, we advocate for the use of block-KSMs for solving Equation 1. Given a matrix $\mathbf{B}\in\mathbb{R}^{d\times m}$ (typically $m\ll d$) with columns $\mathbf{b}^{(1)},\ldots,\mathbf{b}^{(m)}$, the block Krylov subspace is
$\mathcal{K}_{t}(\mathbf{A},\mathbf{B}):=\mathcal{K}_{t}(\mathbf{A},\mathbf{b}^{(1)})+\cdots+\mathcal{K}_{t}(\mathbf{A},\mathbf{b}^{(m)}).$ (14)
That is, $\mathcal{K}_{t}(\mathbf{A},\mathbf{B})$ is the space consisting of all linear combinations of vectors in $\mathcal{K}_{t}(\mathbf{A},\mathbf{b}^{(1)}),\ldots,\mathcal{K}_{t}(\mathbf{A},\mathbf{b}^{(m)})$.
This naturally gives rise to the block-CG algorithm (O’Leary, 1980).
@@
$\mathsf{bcg}_{t}^{(i)}(\mu):=\operatornamewithlimits{argmin}_{\mathbf{x}\in\mathcal{K}_{t}(\mathbf{A},\mathbf{B})}\|\mathbf{A}_{\mu}^{-1}\mathbf{b}^{(i)}-\mathbf{x}\|_{\mathbf{A}_{\mu}}.$ (15)
The block-CG iterates $\mathsf{bcg}_{t}^{(1)}(\mu),\ldots,\mathsf{bcg}_{t}^{(m)}(\mu)$ can be simultaneously computed using $t-1$ block matrix-vector products with $\mathbf{A}$. We discuss implementation and costs further in Section 3.2.
3.1 Main theoretical results
Our first main result is the observation that by augmenting $\mathbf{b}$ with $\mathbf{\Omega}$, block-CG implicitly enjoys the benefits of certain classes of preconditioners built using $\mathbf{\Omega}$. In particular, we have the following error guarantee:
@@ 3.1 Main theoretical results
Fix any matrix $\mathbf{\Omega}\in\mathbb{R}^{d\times m}$ and let $\mathbf{P}=(\mathbf{I}+\mathbf{X})^{-1}$ be any preconditioner where $\operatorname{range}(\mathbf{X})\subseteq\mathcal{K}_{s+1}(\mathbf{A},\mathbf{\Omega})$. Define the augmented starting block $\mathbf{B}=[\mathbf{b}~{}\mathbf{\Omega}]$. Then, for any $t\geq s$, the $t$-th block-CG iterate is related to the $(t-s)$-th preconditioned-CG iterate corresponding to the preconditioner $\mathbf{P}_{\mu}$ in that
@@
In particular, when $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}=\mathbf{A}\langle\mathbf{K}_{s}\rangle$, then the deflation preconditioner $\mathbf{P}_{\mu}$ defined in Equation 8 has the form
$\mathbf{P}^{-1}=\mathbf{I}+\mathbf{X},\text{where $\operatorname{range}(\mathbf{X})\subseteq\mathcal{K}_{s}(\mathbf{A},\mathbf{\Omega})$}.$ (16)
@@
Theorem 3.2 asserts that augmented block-CG automatically performs no worse than Nyström PCG (with the best choice of $s$ and $\theta$) after the same number of matrix-loads. [^4] Recall, $s$ matrix-loads are used to compute the Nyström approximation $\mathbf{A}\langle\mathbf{K}_{s}\rangle$, whose range lives in $\mathcal{K}_{s+1}(\mathbf{A},\mathbf{\Omega})$. This explains the performance observed in Figure 1.
We use Theorem 3.2, and a new bound for Nyström preconditioning (Theorem 4.4), to prove a bound for block-CG reminiscent of Corollary 2.4 for the spectral deflation preconditioner.
Corollary 3.4.
Let $\mathbf{\Omega}\in\mathbb{R}^{d\times(r+2)q}$, where $q\geq\log(1/\delta)/\log(100)$, be a random Gaussian matrix and define the augmented starting block $\mathbf{B}=[\mathbf{b}~{}\mathbf{\Omega}]$. Let
$\varepsilon_{t}(\mu):=2\exp\left(-\frac{t-(3+\log(d))}{3\sqrt{(\lambda_{r+1}+\mu)/(\lambda_{d}+\mu)}}\right).$
Then the block-CG satisfies, with probability at least $1-\delta$,
$\Bigg{\{}\forall\mu\geq 0:\frac{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{bcg}_{t}^{(1)}(\mu)\|_{\mathbf{A}_{\mu}}}{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}\|_{\mathbf{A}_{\mu}}}\leq\varepsilon_{t}(\mu)\Bigg{\}}.$
In particular, after a small burn-in period of $O(\log(d))$ iterations, we have exponential convergence at a rate roughly $\sqrt{(\lambda_{r+1}+\mu)/(\lambda_{d}+\mu)}$. Corollary 3.4 is a special case of a more general bound Theorem 4.5 which takes into account the eigenvalue decay of $\mathbf{A}$.
Remark 3.5.
The fact that Corollary 3.4 guarantees an accurate result for all $\mu\geq 0$ with high probability (as opposed to a single value of $\mu$) will be important in our applications. In particular, it allows guarantees for computing the whole ridge-regression regularization path and will be necessary in the analysis of our algorithm for sampling Gaussian vectors.
Remark 3.6.
Bounds based on the condition number (such as Corollary 3.4) are often pessimistic in practice. To determine how many iterations to run, it is more common to use some sort of a posteriori error estimate. For example, observing the residual $\|\mathbf{b}-\mathbf{A}_{\mu}\mathsf{bcg}_{t}^{(1)}(\mu)\|$ gives some indication of the quality of the solution. More advanced techniques can also be used (Meurant & Tichý, 2024).
@@ 3.2 Computational costs
The block-Lanczos algorithm (which we describe explicitly in Appendix A) applied to $(\mathbf{A},\mathbf{B})$ for $t$ iterations produces $\mathbf{Q}_{t}$ with orthonormal columns and $\mathbf{T}_{t}$ with bandwidth $12m+$ 1 satisfying
$\operatorname{range}(\mathbf{Q}_{t})=\mathcal{K}_{t}(\mathbf{A},\mathbf{B}),\quad\mathbf{Q}_{t}^{\mathsf{T}}\mathbf{A}\mathbf{Q}_{t}=\mathbf{T}_{t}.$ (17)
A standard computation (see Appendix B) reveals that
$\mathsf{bcg}_{k}^{(1)}(\mu)=\|\mathbf{b}\|\mathbf{Q}_{t}(\mathbf{T}_{t}+\mu\mathbf{I})^{-1}\mathbf{e}_{1},$ (18)
where $\mathbf{e}_{1}=[1,0,\ldots,0]^{\mathsf{T}}$. In particular, since $\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{B})=\mathcal{K}_{t}(\mathbf{A},\mathbf{B})$ for any scalar $\mu$, Equation 18 can be used to compute the block-CG iterate for multiple values of $\mu$ using the same $\mathbf{Q}_{t}$ and $\mathbf{T}_{t}$.
@@
Suppose $\mathbf{B}\in\mathbb{R}^{d\times m}$. The block-Lanczos algorithm applied to $(\mathbf{A},\mathbf{B})$ for $t$ iterations produces $\mathbf{Q}_{t}$ and $\mathbf{T}_{t}$ using:
  • $t-1$ matrix-loads of $\mathbf{A}$ (for a total of $m(t-1)$ total matrix-products)
  • $O(dm^{2}t^{2})$ floating point operations operations, and

### v2

@@ 2.2.1 Exact deflation
Traditionally, it has been suggested to take $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ as $\mathbf{A}]\!]_{r}$, the best rank-$r$ approximation to $\mathbf{A}$; i.e. to set $\mathbf{D}$ as the top $r$ eigenvalues of $\mathbf{A}$ and $\mathbf{U}$ the corresponding eigenvectors. In this case Eq. 9 is sometimes called a spectral deflation preconditioner. The name arises because the eigenvalues of $\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}$ are
(11) $\underbrace{\theta+\mu,\ldots,\theta+\mu}_{\text{$r$-copies}},\lambda_{r+1}+\mu,\ldots,\lambda_{d}+\mu,$
and hence, if $\theta\in[\lambda_{d},\lambda_{r+1}]$, then the condition number of $\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}$ is $\kappa_{r+1}(\mu)=(\lambda_{r+1}+\mu)/(\lambda_{d}+\mu)$. In other words, the top $r$ eigenvalues of $\mathbf{A}_{\mu}$ are deflated. In combination with Corollary 2.3, this yields the following convergence guarantee.
@@
Let $\mathbf{P}_{\mu}$ be the preconditioner Eq. 9 corresponding to $\mathbf{A}]\!]_{r}$, the rank-$r$ truncated SVD of $\mathbf{A}$, and $\theta\in[\lambda_{d},\lambda_{r+1}]$. Then the $t$-th PCG iterate (Definition 2.1) corresponding to the preconditioner $\mathbf{P}_{\mu}$ satisfies
$\frac{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{pcg}_{t}(\mu)\|_{\mathbf{A}_{\mu}}}{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}\|_{\mathbf{A}_{\mu}}}\leq 2\exp\left(-\frac{2t}{\sqrt{\kappa_{r+1}(\mu)}}\right).$
When $\lambda_{r+1}\ll\lambda_{1}$ and $\mu$ is not too large, then the rate of convergence guaranteed by the bound can be much faster than without preconditioning (r=0r=0). [^2] Standard CG also satisfies bounds in terms of $\lambda_{r+1}/\lambda_{d}$, at least in exact arithmetic; see Theorem B.1 and the discussion in Appendix B. Of course, any potential benefits to convergence must be weighted against the cost of constructing the spectral deflation preconditioner, and exact deflation, which requires computing the top eigenvectors of $\mathbf{A}$, can be costly.
@@ 2.2.2 Nyström preconditioning
Techniques from randomized numerical linear algebra allow near-optimal low-rank approximations to $\mathbf{A}$ to be computed very efficiently [34, 61]. Intuitively, we might hope that the corresponding preconditioner works nearly as well as spectral deflation.
In particular, it’s reasonable to take $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ as the eigendecomposition of the randomized block-Krylov Nyström approximation, mathematically defined as
(12) $\mathbf{A}\langle\mathbf{K}_{s}\rangle:=(\mathbf{A}\mathbf{K}_{s})(\mathbf{K}_{s}^{\mathsf{T}}\mathbf{A}\mathbf{K}_{s})^{\dagger}(\mathbf{K}_{s}^{\mathsf{T}}\mathbf{A}),$
where $\mathbf{K}_{s}:=[\mathbf{\Omega}~\mathbf{A}\mathbf{\Omega}~\cdots~\mathbf{A}^{s-1}\mathbf{\Omega}]\in\mathbb{R}^{d\times(s\ell)}$ and $\mathbf{\Omega}\in\mathbb{R}^{d\times\ell}$ is a matrix of independent standard normal random variables. [^3] In this paper, we only consider Nyström preconditioning where $\mathbf{\Omega}$ is a Gaussian matrix. Variants of Nyström approximation based on subsampling rows/columns of $\mathbf{A}$ are effective in some settings (e.g. Kernel Ridge Regression); see [23, §2.2.3] for a discussion. This variant of the Nyström approximation is among the most powerful randomized low-rank approximation algorithms, and can be implemented using $s$ matrix-loads. Note, however, that the algorithm should not be implemented as written in Eq. 12. In particular, one should avoid using a monomial basis $\mathbf{K}_{s}$ and carefully structure the interactions with $\mathbf{A}$ to avoid unnecessary costs; see e.g. [61] for pseudocode.
Nyström-based preconditioning is effective in theory and practice, and has been an active area of research in recent years [40, 23, 5, 20, 65, 36, 15]. Most related to the content of this paper is the theoretical analysis of [23], which proves that if $s=1$, $\theta=\lambda_{\ell}(\mathbf{A}\langle\mathbf{\Omega}\rangle)$, and the sketching dimension $\ell$ is on the order of the effective dimension $d_{\textup{eff}}(\mu):=\operatorname{tr}\bigl{(}\mathbf{A}\mathbf{A}_{\mu}^{-1}\bigr{)}=\sum_{i=1}^{d}{\lambda_{i}}/{(\lambda_{i}+\mu)}$, then Nyström PCG converges in at most $t=O(\log(1/\varepsilon))$ iterations; i.e. independent of any spectral properties of $\mathbf{A}$. Our analysis makes use of the same general techniques as [23], but is applicable when $s>1$ as well as if $\mu=0$. We discuss the theoretical bounds of [23] in Section B.2.
@@ 3 Our approach: augmented block-CG
In this paper, we advocate for the direct application of block KSMs to Eq. 1. The concept of block Krylov subspaces Eq. 6 naturally gives rise to the block-CG algorithm [48].
@@
$\mathsf{bcg}_{t}^{(i)}(\mu):=\operatornamewithlimits{argmin}_{\mathbf{x}\in\mathcal{K}_{t}(\mathbf{A},\mathbf{B})}\|\mathbf{A}_{\mu}^{-1}\mathbf{b}^{(i)}-\mathbf{x}\|_{\mathbf{A}_{\mu}}.$
The block-CG iterates $\mathsf{bcg}_{t}^{(1)}(\mu),\ldots,\mathsf{bcg}_{t}^{(m)}(\mu)$ can be simultaneously computed using t1t-1 block matrix-vector products with $\mathbf{A}$. We discuss implementation and costs further in Section 3.2.
3.1 Implicit preconditioning
Our main conceptual result is that by augmenting $\mathbf{b}$ with $\mathbf{\Omega}$, block-CG implicitly enjoys the benefits of certain classes of preconditioners built using $\mathbf{\Omega}$. We begin with a key observation about the relation between the block Krylov subspace Eq. 6 and the preconditioned Krylov subspace Eq. 8 corresponding to a certain class of preconditioners.
@@ 3.1 Implicit preconditioning
Suppose $\mathbf{P}_{\mu}=(\mathbf{I}+\mathbf{X})^{-1}$, where $\operatorname{range}(\mathbf{X})\subseteq\mathcal{K}_{s+1}(\mathbf{A}_{\mu},\mathbf{\Omega})$. Then,
The basic idea is simple. By definition, $\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{b};\mathbf{P}_{\mu})$ consists of linear combinations of the vectors
(13) $(\mathbf{P}_{\mu}^{-1}\mathbf{A})^{k}\mathbf{P}_{\mu}^{-1}\mathbf{b}=((\mathbf{I}+\mathbf{X})\mathbf{A})^{k}(\mathbf{I}+\mathbf{X})\mathbf{b},\quad k=0,1,\ldots,t-1,$
and each $((\mathbf{I}+\mathbf{X})\mathbf{A})^{k}(\mathbf{I}+\mathbf{X})\mathbf{b}$ can be expressed as linear combination of vectors which live in the specified space.
Proof 3.3.
(14) $\mathcal{K}_{1}(\mathbf{P}^{-1}\mathbf{A},\mathbf{P}^{-1}\mathbf{b})=\operatorname{span}(\mathbf{P}^{-1}\mathbf{b})=\operatorname{span}((\mathbf{I}+\mathbf{X})\mathbf{b})=\operatorname{span}(\mathbf{b}+\mathbf{X}\mathbf{b}).$
Now, assume that
(15) $\mathcal{K}_{t-1}(\mathbf{P}^{-1}\mathbf{A},\mathbf{P}^{-1}\mathbf{b})\subseteq\mathcal{K}_{t-1}(\mathbf{A},\mathbf{b})+\mathcal{K}_{(t-1)+s}(\mathbf{A},\mathbf{\Omega}).$
We consider the order $t$ subspace
(16) $\hskip 130.0002pt(\mathbf{P}^{-1}\mathbf{A})^{t-2}\mathbf{P}^{-1}\mathbf{b},(\mathbf{P}^{-1}\mathbf{A})^{t-1}\mathbf{P}^{-1}\mathbf{b}\}.$
From the inductive hypothesis, we know that for all j=0,,t2j=0,...,t-2,
(17) $(\mathbf{P}^{-1}\mathbf{A})^{j}\mathbf{P}^{-1}\mathbf{b}$ $\in\mathcal{K}_{t-1}(\mathbf{P}^{-1}\mathbf{A},\mathbf{P}^{-1}\mathbf{b})$
(18) $\subseteq\mathcal{K}_{t-1}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s-1}(\mathbf{A},\mathbf{\Omega})$
(19) $\subseteq\mathcal{K}_{t}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s}(\mathbf{A},\mathbf{\Omega}),$
(20) $(\mathbf{P}^{-1}\mathbf{A})^{t-1}\mathbf{P}^{-1}\mathbf{b}=(\mathbf{P}^{-1}\mathbf{A})\mathbf{v}=(\mathbf{I}+\mathbf{X})\mathbf{A}\mathbf{v}=\mathbf{A}\mathbf{v}+\mathbf{X}\mathbf{A}\mathbf{v}.$
Clearly $\mathbf{X}\mathbf{A}\mathbf{v}\subset\operatorname{range}(\mathbf{X})\subseteq\mathbf{K}_{s+1}(\mathbf{A},\mathbf{\Omega})$. Moreover, as noted, $\mathbf{v}\in\mathcal{K}_{t-1}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s-1}(\mathbf{A},\mathbf{\Omega})$ and so
(21) $\mathbf{A}\mathbf{v}\in\mathbf{A}\mathcal{K}_{t-1}(\mathbf{A},\mathbf{b})+\mathbf{A}\mathcal{K}_{t+s-1}(\mathbf{A},\mathbf{\Omega})\subseteq\mathcal{K}_{t}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s}(\mathbf{A},\mathbf{\Omega}).$
This proves the result.
Theorem 3.2 allows the performance of augmented block-CG to be related to PCG with a broad choice of preconditioner.
Theorem 3.4.
Fix any matrix $\mathbf{\Omega}\in\mathbb{R}^{d\times m}$ and let $\mathbf{P}=(\mathbf{I}+\mathbf{X})^{-1}$ be any preconditioner where $\operatorname{range}(\mathbf{X})\subseteq\mathcal{K}_{s+1}(\mathbf{A},\mathbf{\Omega})$. Define the augmented starting block $\mathbf{B}=[\mathbf{b}~\mathbf{\Omega}]$. Then, for any $t\geq s$, the $t$-th block-CG iterate is related to the (ts)(t-s)-th preconditioned-CG iterate corresponding to the preconditioner $\mathbf{P}_{\mu}$ in that
@@
Proof 3.5.
Clearly $\mathcal{K}_{t}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s}(\mathbf{A},\mathbf{\Omega})\subseteq\mathcal{K}_{t+s}(\mathbf{A},[\mathbf{b}~\mathbf{\Omega}])$, so the result follows immediately from Theorem 3.2, the optimality of block-CG, and the definition of preconditioned-CG.
@@
Theorem 3.4 asserts that augmented block-CG automatically performs no worse than Nyström PCG (with the best choice of $s$ and $\theta$) after the same number of matrix-loads. [^4] Recall, $s$ matrix-loads are used to compute the Nyström approximation $\mathbf{A}\langle\mathbf{K}_{s}\rangle$, whose range lives in $\mathcal{K}_{s+1}(\mathbf{A},\mathbf{\Omega})$. In particular, when $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}=\mathbf{A}\langle\mathbf{K}_{s}\rangle$, then the deflation preconditioner $\mathbf{P}_{\mu}$ defined in Eq. 9 has the form
(22) $\mathbf{P}^{-1}=\mathbf{I}+\mathbf{X},\quad\text{where $\operatorname{range}(\mathbf{X})\subseteq\mathcal{K}_{s+1}(\mathbf{A},\mathbf{\Omega})$}.$
This explains the performance observed in Fig. 1.
In Section 4, we use use Theorem 3.4, and a new bound for Nyström preconditioning (Theorem 4.7), to prove a bound for block-CG reminiscent of Corollary 2.4 for the spectral deflation preconditioner.
@@ 3.2 Computational costs
The block-Lanczos algorithm applied to $(\mathbf{A},\mathbf{B})$ for $t$ iterations produces $\mathbf{Q}_{t}$ with orthonormal columns and $\mathbf{T}_{t}$ with bandwidth $2m+1$ satisfying
(23) $\operatorname{range}(\mathbf{Q}_{t})=\mathcal{K}_{t}(\mathbf{A},\mathbf{B}),\quad\mathbf{T}_{t}=\mathbf{Q}_{t}^{\mathsf{T}}\mathbf{A}\mathbf{Q}_{t}.$
A standard computation reveals that
(24) $\mathsf{bcg}_{k}^{(1)}(\mu)=\|\mathbf{b}\|\mathbf{Q}_{t}(\mathbf{T}_{t}+\mu\mathbf{I})^{-1}\mathbf{e}_{1},$
where $\mathbf{e}_{1}=[1,0,\ldots,0]^{\mathsf{T}}$. In particular, since $\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{B})=\mathcal{K}_{t}(\mathbf{A},\mathbf{B})$ for any scalar $\mu$, Eq. 24 can be used to compute the block-CG iterate for multiple values of $\mu$ using the same $\mathbf{Q}_{t}$ and $\mathbf{T}_{t}$. For more details on block-Lanczos and the connection to block-CG, see [48, 55, 7, 60, etc.].
There are many variants of block-Lanczos, since there are many possible orthogonlization schemes that can be use within block-Lanczos type algorithms [6, 47, 1]. The following provides a high-level overview of the costs associated with block-Lanczos with full-reorthogonalization.
@@
Suppose $\mathbf{B}\in\mathbb{R}^{d\times m}$. The block-Lanczos algorithm (with full orthogonalization) applied to $(\mathbf{A},\mathbf{B})$ for $t$ iterations produces $\mathbf{Q}_{t}$ and $\mathbf{T}_{t}$ using:
  • t1t-1 matrix-loads of $\mathbf{A}$ (for a total of m(t1)m(t-1) total matrix-products)
  • $O(dm^{2}t^{2})$ floating point operations, and
@@
Proof 3.8 (Proof of Theorem 3.7).
At each iteration the block-Lanczos algorithm performs one matrix-load ($m$ parallel matrix products) and lower-order arithmetic. The dominant storage cost is storing $\mathbf{Q}_{t}$, which has $mt$ columns. The arithmetic cost is dominated by the $O(dm^{2}t^{2})$ operations required to obtain $\mathbf{Q}_{t}$ using full reorthogonalization. Since $\mathbf{T}_{t}$ (and hence $\mathbf{T}_{t}+\mu\mathbf{I}$) are $(mt)\times(mt)$ matrices of bandwidth $O(m)$, the linear system $(\mathbf{T}_{t}+\mu\mathbf{I})\mathbf{z}=\mathbf{e}_{1}$ can be solved in $O((mt)m^{2})=O(m^{3}t)=O(dm^{2})$ time. Subsequently, since $\mathbf{Q}_{t}$ is a $d\times(mt)$ matrix, $\mathbf{Q}_{t}\mathbf{z}$ can be computed in $O(dmt)$ time.